what is the magnitude of the electrostatic force between two electrons separated by a distance of 1.00 X 10-8 meter?


1
Expert's answer
2021-11-02T10:00:59-0400

The magnitude of the electrostatic force between two electrons is given as follows:


"F = k\dfrac{e^2}{r^2}"

where "k = 8.99\times 10^9Nm^2/C^2" is the Coulomb constant, "e = 1.6\times 10^{-19}C" is the charge of electron, "r = 10^{-8}m" is the distance between electrons. Thus, obtain:


"F = 8.99\times 10^9\dfrac{(1.6\times 10^{-19})^2}{(10^{-8})^2} \approx 2.30\times 10^{-12}N"

Answer. "2.30\times 10^{-12}N".